# Accelerating Materials Design via LLM-Guided Evolutionary Search

### v1

@@ D.2 Additional Details
5. Motif replacement: Replace a structural fragment with another serving a similar role.
$\text{ABC-ring}\;\;\rightarrow\;\;\text{DEF-ring}$
@@
7. Layered intercalation: Insert atoms between layers in a layered structure.
$\text{[ABC]}\;\;\rightarrow\;\;\text{[ABC]}\cdot\text{D}$
@@
10. Structural isomer generation: Rearrange atomic connectivity while preserving formula.
$\text{A{-}B{-}C{-}D}\;\;\rightarrow\;\;\text{A{-}C{-}B{-}D}$
11. Group-based recombination: Merge fragments from two known compounds.
$\text{(A{-}B{-}C) + (D{-}E{-}F)}\;\;\rightarrow\;\;\text{A{-}E{-}C}$
@@
15. Phase diagram extrapolation: Propose compounds between two known stable ones.
$\text{(A{-}B), (B{-}C)}\;\;\rightarrow\;\;\text{A{-}C}\quad\text{or}\quad\text{A}_{0.5}\text{B}_{0.2}\text{C}_{0.3}$
@@ E.1 Case Study
To better understand how LLEMA’s evolutionary mechanism balances exploration and exploitation, we analyze the temporal dynamics of the search process across iterations. This case study focuses on three aspects of the evolutionary trajectory: (i) memorization rate the proportion of generated compounds retrieved directly from known databases such as the Materials Project; (ii) chemical diversity the spread of elemental coverage across the periodic table; and (iii) validity progression the fraction of syntactically and physically valid compounds discovered over successive generations.
@@
From Memorization to Exploration.
In the early stages of optimization, LLEMA’s search behavior is dominated by memorization, with a substantial portion of generated candidates overlapping with known entries from the Materials Project (Figure 8). As the search evolves, the proportion of externally sourced structures rapidly declines. From about 83% initially, the overlap with Materials Project drops to 10% by 250 iterations, eventually dropping to about 3% by the end of the evolution. This steady decrease highlights a clear transition from memorization of existing chemical knowledge to exploration of novel compositions. By later iterations, the candidate pool is largely composed of previously unseen structures, indicating that the model has shifted from recall-driven synthesis toward genuine materials discovery.
@@
Tracking candidates in property space reveals how LLEMA’s population progressively migrates toward feasible and optimal regions over time. As shown in Figure 9, at early stages stages (Iteration 250), only about 30% of SAW/BAW acoustic substrates satisfy physical validity constraints, with most scattering across thermodynamically unstable or suboptimal zones. As the evolution proceeds, this fraction increases to nearly 46% by iteration 1000, reflecting the growing influence of rule-based chemical filters and adaptive feedback mechanisms. Concurrently, the Pareto front advances steadily, expanding the achievable trade-off frontier and uncovering materials that balance multiple performance criteria. This improvement stems from LLEMA’s evolutionary refinement process, where chemically guided mutations, crossover between promising candidates, and memory-based selection pressure iteratively prune the search space. The result is a guided transition from memorized priors to data-driven innovation, leading to denser clusters of thermodynamically stable, property-aligned materials and a measurable improvement in both diversity and discovery quality across generations.
@@
As memorization declines, the fraction of valid and physically plausible candidates rises steadily. Early populations contain roughly 30% valid materials—many of which are simple substitutions of known prototypes—while later generations surpass 80% validity. This improvement results from the combined effect of oracle-guided scoring and constraint-aware feedback that prune chemically inconsistent structures while reinforcing successful design patterns. The process simultaneously expands the search’s coverage of the periodic table: early iterations are dominated by light elements and common oxides, whereas later generations incorporate diverse transition metals, alkaline earths, and rare-earth substitutions, reflecting a richer exploration of the underlying chemical landscape.
E.2 Diversity of Identified Materials

### v2

@@ B.4 Evolutionary Generation Rules
Appendix C Datasets and Benchmark
To evaluate multi-objective material discovery, we curate a diverse benchmark spanning 14 representative design tasks across various domains (see Table 2). Each task defines a distinct combination of physicochemical constraints that reflect practical design objectives in real-world materials engineering. Together, the benchmark captures the breadth of challenges faced in materials design, from optimizing performance–stability trade-offs to balancing mechanical, dielectric, and sustainability objectives. The resulting dataset assesses how well models can reason over complex, interdependent physical properties and generate synthesizable candidates under realistic constraints.
Wide-Bandgap Semiconductors.
SAW/BAW Acoustic Substrates.
High-$k$ Dielectrics.
Solid-State Electrolytes.
Solid electrolytes promise safe, high-energy batteries by replacing flammable liquid electrolytes. Candidates must be thermodynamically stable (formation energy $\leq-1.0$ eV/atom), electronically insulating (band gap $\geq 2.0$ eV), and contain mobile species such as Li, Na, K, Mg, Ca, or Al. This task reflects the fundamental trade-off between chemical stability and ionic conductivity, which is central to enabling next-generation solid-state batteries.
Piezo Energy Harvesters.
Transparent Conductors.
Electrically Insulating Dielectrics.
Photovoltaic Absorbers.
Photovoltaic materials must absorb sunlight efficiently while remaining stable, earth-abundant, and non-toxic. Target absorbers have optimal band gaps (0.70.72.02.0 eV) for solar conversion, formation energies $\leq 0.0$ eV/atom, and must exclude rare or hazardous elements. This task reflects real-world sustainability constraints, forcing models to move beyond theoretical optima and identify candidates suitable for large-scale, affordable solar deployment.
Hard Coating Materials.
Hard, Stiff Ceramics.
Structural Materials for Aerospace.
Acousto-Optic Hybrids.
Materials for acousto-optic devices must balance piezoelectric and dielectric properties to minimize loading while enabling efficient coupling. Candidates are required to exhibit piezoelectric coefficients in the range $3\leq d_{ij}\leq 9$ pC/N and dielectric constants in the range $8\leq\kappa\leq 85$, with a preference for a narrow $\kappa$ band to reduce dielectric loading. Ranking emphasizes proximity to target $d$-bands and mid-range $\kappa$ values, highlighting the trade-off between acoustic response and dielectric stability.
Low-Density Structural Materials.
Toxic-Free Perovskite Oxides.
Appendix D Implementation Details
D.1 Baselines
CDVAE.
CDVAE (Xie et al., 2022) is a conditional variational autoencoder tailored for crystal structure generation. It learns latent representations of crystals conditioned on composition, enabling the generation of valid and diverse candidate structures. We implement CDVAE using the official open-source repository [^5] https://github.com/txie-93/cdvae with default parameters, outputting the generated structures in CIF (Crystallographic Information File) format, along with latent embeddings stored as NumPy arrays for downstream analysis.
G-SchNet.
G-SchNet (Gebauer et al., 2019) is a graph-based deep generative model for molecular and crystal structures, built on the SchNet architecture. It incrementally generates atom types and positions conditioned on the partially built structure, allowing it to capture geometric and chemical validity. We implement G-SchNet using the open-source codebase [^6] https://github.com/atomistic-machine-learning/G-SchNet with default hyperparameters for $15000$ iterations. The generated samples are stored in a database, which is subsequently converted into candidate-specific CIF files.
DiffCSP.
MatterGen.
We use the official mattergen [^8] https://github.com/microsoft/mattergen release and checkpoints. For candidate generation, we sample batches from the released base checkpoint. Generated candidates are then evaluated using surrogate predictors and the Materials Project API to determine the candidate quality.
End2end.
We implemented a pipeline with base LLM as GPT-4o-mini to generate candidates conditioned on the design task and its corresponding property constraints. The LLM operated with a sampling temperature of $\tau=0.8$ to promote diversity while preserving structural coherence. It was explicitly instructed to output candidates in Crystallographic Information File (CIF) format, ensuring standardized structural representations suitable for subsequent validation and property evaluation using the surrogate-assisted oracle prediction.
LLMatDesign.
LLMatDesign (Jia et al., 2024) leverages LLMs for material design by prompting LLMs to iteratively improve the provided material to return the material with the single target property. As LLMatDesign is primarily designed for single-objective optimization, we adapt its released implementation [^9] https://github.com/Fung-Lab/LLMatDesign for our multi-objective materials discovery benchmarks and add surrogate-augmented oracle from our pipeline for feedback. For candidate generation, we adopt their original prompt template, which provides a material’s chemical formula and task-specific properties and asks the model to propose a modification that satisfies the constraints. The model selects one of four modification types among ‘exchange’, ‘substitute’, ‘remove’, or ‘add’, and outputs both the modification and a natural-language hypothesis justifying the change.
D.2 LLEMA
Material Design.
Figure 6 illustrates an example prompt for the Wide-Bandgap Semiconductor task. The prompt begins with general instructions specifying the LLM’s role and objective, followed by task-specific details such as the description of the design problem, property constraints (e.g., required band gap and formation energy), and a set of previously explored candidates with their associated property values sampled from the experience buffer. We then sample $b=2$ candidate outputs at a temperature of $\tau=0.8$, chosen to balance creativity with adherence to constraints while exploiting promising directions in the search space. To further ensure physical plausibility, we introduce rule-based constraints during the evolutionary phase (after the initial generation step, $n=0$), which guide the LLM through the material discovery cycle and promote chemically valid candidates. We randomly sample a subset of 6 rules from the set of rules and provide it to the LLM to guide the candidate generation. By providing task descriptions, prior evaluations, and rule-based constraints in the prompt, we effectively steer the LLM toward feasible, property-aligned material candidates.
Crystallographic Representation.
Data-Driven Evaluation.
Figure 6: Example of input prompts for the wide-bandgap semiconductors task. (a) Candidate Generation Step , including task specification, evolution constraints, in-context demonstrations, and (b) Crystallographic File (CIF) Representation Step containing the CIF generation prompt.
Experience Management.
We adopt an island-based evolutionary strategy (Romera-Paredes et al., 2024; Shojaee et al., 2025a; Abhyankar et al., 2025) to manage the experience buffer, where generated material candidates and their evaluation scores are distributed across $m=5$ independently evolving islands. Each island is initialized with a small set of seed candidates sampled from the Materials Project API and refined using the LLM. Within each island, the buffer is divided into two components: a reward buffer, which stores candidates that fully satisfy all task-specific constraints (e.g., band gap and formation energy), and an error buffer, which stores candidates that only partially satisfy or completely violate the constraints. This separation enables the framework to reinforce high-quality generations while also retaining failure cases, which serve as negative examples to guide exploration. The experience buffer is further leveraged to construct prompts for subsequent LLM calls. After the prompt template is updated with task-specific information, one of the $m$ islands is selected, and $k=2$ candidates are sampled from its buffers to serve as in-context demonstrations. Candidate selection follows a Boltzmann strategy (De La Maza and Tidor, 1992) that assigns higher probability to clusters with stronger evaluation scores. Specifically, if $s_{i}$ denotes the score of the $i$-th cluster, the probability $P_{i}$ of selecting it is given by:
$P_{i}=\frac{\exp(s_{i}/\tau_{c})}{\sum_{i}\exp(s_{i}/\tau_{c})},\qquad\tau_{c}=T_{0}\left(1-\frac{u\bmod M}{M}\right),$
where $\tau_{c}$ is the temperature parameter, $u$ is the current number of candidates on the island, and $T_{0}=0.1$ and M=10, [^0] M=10{,}000 are hyperparameters. Once a cluster is selected, we sample candidates from it for inclusion in the next generation. This mechanism integrates information from both successful and failed generations while preserving diversity across islands, thereby guiding the LLM toward more effective material discovery.
@@ E.1 Case Study
To better understand how LLEMA’s evolutionary mechanism balances exploration and exploitation, we analyze the temporal dynamics of the search process across iterations. This case study focuses on three aspects of the evolutionary trajectory: (i) memorization rate: the proportion of generated compounds retrieved directly from known databases such as the Materials Project; (ii) chemical diversity: the spread of elemental coverage across the periodic table; and (iii) validity progression: the fraction of syntactically and physically valid compounds discovered over successive generations.
From Memorization to Exploration.
In the early stages of optimization, LLEMA’s search behavior is dominated by memorization, with a substantial portion of generated candidates overlapping with known entries from the Materials Project (Figure 7). As the search evolves, the proportion of externally sourced structures rapidly declines. From about 83% initially, the overlap with Materials Project drops to 10% by 250 iterations, eventually dropping to about 3% by the end of the evolution. This steady decrease highlights a clear transition from memorization of existing chemical knowledge to exploration of novel compositions. By later iterations, the candidate pool is largely composed of previously unseen structures, indicating that the model has shifted from recall-driven synthesis toward genuine materials discovery. Figure 8 further highlights how the evolution process guides the exploration toward a more diverse set of elements.
@@
Tracking candidates in property space reveals how LLEMA’s population progressively migrates toward feasible and optimal regions over time. As shown in Figure 9, at early stages (Iteration 250), only about 30% of SAW/BAW acoustic substrates satisfy physical validity constraints, with most scattering across thermodynamically unstable or suboptimal zones. As the evolution proceeds, this fraction increases to nearly 46% by iteration 1000, reflecting the growing influence of rule-based chemical filters and adaptive feedback mechanisms. Concurrently, the Pareto front advances steadily, expanding the achievable trade-off frontier and uncovering materials that balance multiple performance criteria. This improvement stems from LLEMA’s evolutionary refinement process, where chemically guided mutations, crossover between promising candidates, and memory-based selection pressure iteratively prune the search space. The result is a guided transition from memorized priors to data-driven innovation, leading to denser clusters of thermodynamically stable, property-aligned materials and a measurable improvement in both diversity and discovery quality across generations.
@@
As memorization declines, the fraction of valid and physically plausible candidates rises steadily. Early populations contain roughly 30% valid materials, many of which are simple substitutions of known prototypes, while later generations surpass 80% validity. This improvement results from the combined effect of oracle-guided scoring and constraint-aware feedback that prunes chemically inconsistent structures while reinforcing successful design patterns. The process simultaneously expands the search’s coverage of the periodic table: early iterations are dominated by light elements and common oxides, whereas later generations incorporate diverse transition metals, alkaline earths, and rare-earth substitutions, reflecting a richer exploration of the underlying chemical landscape.
E.2 Robustness to Prompt Variations
LLM-based generation can be sensitive to how prompts are phrased, so we assess whether LLEMA remains stable under changes in prompt structure. Starting from the base LLEMA prompt, we manually created three alternative versions by rephrasing instructions, reorganizing constraint descriptions, and varying the emphasis on key design objectives. These variants preserve the same semantics while altering the surface form of the prompt. We then evaluate whether LLEMA continues to produce consistent and chemically plausible candidates across these prompt variations. To assess robustness, we performed a focused study on two representative tasks: Wide-Bandgap Semiconductors and Hard, Stiff Ceramics. For each task, we executed the design iterations while varying the prompt formulations used during generation. Across these controlled experiments, we observed only modest variation in the number of valid candidates produced, indicating that the method is reasonably stable under moderate prompt modifications. Results are summarized in Table 7.
Table 7: Effect of different prompt formulations on valid candidate generation (hit-rate).
Prompt Version
Wide-bandgap
Semiconductors
Hard, stiff
Ceramics
Version 1 24.0 41.3
Version 2 22.8 42.8
Version 3 26.8 44.5
E.3 Effect of Model Scale and Domain Knowledge
Our primary experiments use GPT-4o-mini and Mistral-Small-3.2, both of which are mid-scale models. We further evaluate open-sourced models like Qwen2.5-7B/14B/32B (Yang et al., 2024a) to function within the pipeline. However, generating valid CIF files and proposing plausible materials is not a simple rule-based task, as LLMs must encode enough materials-science domain knowledge to respect crystallographic constraints, chemical compatibility, and physically realistic structures. This explains why performance improves from 7B to 14B/32B with the model size primarily acting as a proxy for embedded scientific priors (see Table 8). This also underscores a central point: rule-based evolution alone is not enough. Without domain knowledge, search trajectories drift toward non-physical or chemically invalid structures even if the output format is correct. Larger models succeed not because of scale, but because they carry a richer implicit understanding of stability, bonding, and symmetry. Thus, the framework leverages the domain knowledge encoded in the LLM, not the rules alone.
Table 8: Hit rates (%) across tasks for different model sizes.
Task Qwen2.5-7B Qwen2.5-14B Qwen2.5-32B
High-$k$ Dielectrics 4.46 10.20 13.27
SAW/BAW Acoustic Substrates 1.30 20.83 23.81
Wide-Bandgap Semiconductors 9.80 22.55 20.00
E.4 Reliability of Surrogate Models
Density Functional Theory (DFT) remains the gold standard for evaluating stability and functional properties; however, its high computational cost (tens of minutes to hours per structure) makes it impractical to incorporate directly into an evolutionary loop. To maintain fast iteration, LLEMA employs surrogate models as lightweight proxies, enabling rapid property estimation while guiding the search toward promising regions of the design space. To assess how well surrogate-guided discoveries translate to high-fidelity physics, we sample 150 valid candidates across four tasks: Wide-Bandgap Semiconductors, Photovoltaic Absorbers, Piezo Energy Harvesters, and Acousto-optic Hybrids, and evaluate them using Quantum ESPRESSO [^10] https://github.com/QEF (Giannozzi et al., 2009) using default parameters. Among these candidates, 144 out of 150 (96%) satisfy the task constraints under DFT, indicating strong agreement between surrogate predictions and high-fidelity outcomes. As shown in Table 9, the fraction of LLEMA-generated candidate materials whose predicted properties are physically consistent with DFT reference calculations is provided. High DFT validity across all five evaluated properties indicates that the surrogate models used within LLEMA reliably approximate true quantum-mechanical behavior, supporting the robustness of property predictions across diverse materials design tasks.
Table 9: DFT Validity of predicted material properties across different tasks.
Property DFT Validity
band_gap 97%
formation_energy 99%
energy_above_hull 96%
piezo_max_dij 97%
piezo_max_dielectric 98%
E.5 Robustness to Noise
Since LLEMA ’s evolutionary loop depends on surrogate model predictions to guide the search, it is important to assess whether small perturbations in these predictions alter the trajectory of the algorithm. To test robustness, we constructed noisy surrogate models by injecting additive Gaussian noise ($\mu$ = 0, $\sigma$ = 0.05) into every predicted property at every evaluation step. The rest of the pipeline, including mutation, selection, and validity checks, was kept identical to the standard LLEMA setup. We then re-ran LLEMA with these noisy surrogates across all tasks and compared the resulting candidate sets with those obtained using the original models. While the overall validity rate decreases, LLEMA consistently identifies chemically similar regions of the design space: we observe 100% elemental overlap for Wide-Bandgap Semiconductors and High-k Dielectrics, 96% for SAW/BAW Acoustics, and 84% for Hard, Stiff Ceramics. This high overlap indicates that LLEMA does not simply follow surrogate feedback, but leverages its domain knowledge and evolutionary reasoning to steer the search toward plausible solutions even when the guidance signal is perturbed. To further assess high-fidelity correctness under noisy guidance, we performed DFT calculations on a set of 150 candidates generated with noisy surrogates. 141 out of 150 (94%) satisfy the task constraints under DFT, closely matching the noiseless evaluation. Together, these results show that LLEMA maintains coherent, physically meaningful trajectories even under uncertain predictions, reinforcing the robustness and reliability of the method.
E.6 Efficiency Analysis
[…]
Table 10: Efficiency comparison across methods.
Method Input Tokens/Call Output Tokens/Call Time/Iter. (s) API Calls/Iter.
LLEMA 563 176 15.84 3
LLMatDesign 509 185 13.23 3
Base LLM 457 118 4.12 2
E.7 Diversity of Identified Materials